# bibliography.bib

### v1

@@ I INTRODUCTION
Unmanned Aerial Vehicles (UAVs) are increasingly employed in civilian applications, including aerial surveillance, precision agriculture, and logistics [shakhatreh2019unmanned]. As critical components of emerging urban air mobility (UAM) systems, UAVs rely on networked communication and onboard autonomy, rendering them vulnerable to sophisticated cyber-attacks [panda2024action]. Adversaries may intercept telemetry data and inject falsified information by gaining unauthorized access to communication channels, thereby obstructing the UAM operator’s ability to verify compliance with planned flight paths as shown in Fig. 1. Failing to detect adversarial attacks as general anomalies can allow malicious actors to bypass security measures and inflict serious harm on airspace operations. This may lead to UAVs breaching restricted or unauthorized zones, thereby endangering other airspace users and heightening the risk of hijacking or unintentional exposure of sensitive information. Therefore, enhancing the security of UAV systems is crucial—especially in dynamic, time-critical operational environments. Traditional security methods such as cryptography and anomaly-based IDS face challenges like latency, high false positives, and poor generalization to novel threats [al2024machine].
@@
Generative adversarial networks (GANs) have emerged as a powerful tool for crafting polymorphic cyberattacks that evade detection. Prior work has demonstrated their use in generating adversarial examples that mimic legitimate data distributions while compromising classifier performance [chauhan2020polymorphic, gu2023gan]. However, distinguishing such stealthy adversarial samples from natural out-of-distribution (OOD) samples remains a major challenge. Adversaries can exploit this weakness by generating perturbations that resemble benign anomalies, thereby bypassing anomaly detection-based IDS. An inability to differentiate out-of-distribution (OOD) samples from adversarial attacks can undermine the effectiveness of conventional filtering methods, such as the extended Kalman filter (EKF) [xiao2022cyber], which may struggle to detect stealthy adversarial inputs due to their statistical resemblance to the assumed OOD features.
Statistical methods for adversarial detection [pang2018towards, smith2018understanding] often assume fixed data distributions and lack adaptability to high-dimensional stealthy attacks. Moreover, they fail to model vulnerabilities in the latent space, limiting their effectiveness in detecting unseen or well-crafted adversarial inputs. To overcome these limitations, deep generative models such as variational autoencoders (VAEs) have been proposed [takiddin2021variational]. VAEs capture probabilistic latent representations that support better anomaly detection. However, their generalization ability is limited in the presence of multiple attack types. Conditional VAEs (CVAEs) [pol2019anomaly], which incorporate attack labels into the learning process, enhance discriminative power and robustness—but their application in detecting and characterizing diverse stealthy attacks remains underexplored. To address these gaps, this paper introduces a novel framework that combines generative adversarial evasion with probabilistic detection using CVAE for UAV cyber-physical systems. The key contributions are summarized as:
@@
  • Probabilistic Detection via CVAE: To detect such stealthy attacks from OOD samples, we implement a CVAE-based detection system. Using negative log-likelihood as a detection score, our method outperforms traditional Mahalanobis-based detectors [pang2018towards] and advanced regret-based CVAE [xiao2020likelihood] baselines in distinguishing stealthy attacks from genuine OOD samples.
@@ II-B Cyber Attack Dataset
The UAV cyber-attack dataset utilized in this study is based on the intrusion detection framework developed by Hassler et al. [hassler2023cyber], where a comprehensive experimental setup was employed. The dataset is comprised of UAV telemetry obtained after several types of cyberattacks, briefly described as follows:
@@
These attacks exhibit clear manifestations across both cyber and physical domains. Regarding the cyber features, only those attributes common across all attack types were retained to ensure consistency. These include: timestamp, frame number, frame length, transmission duration over the wireless medium, frame sequence number, control type and subtype, and the WLAN fragment number. In contrast, the physical features encompass various UAV telemetry data, such as altitude; velocities along the $x,y,z$ axes pitch, roll, yaw; flight distance and time; pressure; battery status; and spatial distances along the $x,y,z$ axes relative to the Tello Pad. Details of the experiment is provided in [hassler2023cyber].
@@ III-B Discriminator
During the cGAN training, we assume that the attacker has the access to the IDS via an external API. The cGAN can be trained in a standard manner as shown in [xu2019modeling].
@@ III-C Iterative Refinement for Stealthy Attack
Iterative refinement strategy is analogous to the one in projected gradient descent (PGD) attacks [madry2018towards], where attack is injected in multi-steps so that the overall perturbations stay within a certain bound. The aim is to make the adversarial attack $X_{\textup{adv}}$ close to the OOD sample $X_{\textup{ood}}$ generated from normally distributed noise with the noise scale $\varrho$, defined as $X_{\text{ood}}=X_{\text{att}}+\mathcal{N}\left(\mathbf{0},{\varrho}^{2}\mathbf{I}\right)$ Hence, to ensure the stealth nature of the adversarial attack, $X_{\text{adv}}$ needs to be distributionally similar to $X_{\text{ood}}$, with high attack success rate. Hence, we define the stealthy adversarial attacks as follows:
@@ IV Conditional Variational Autoencoder for Stealthy Attack Detection
CVAEs [pol2019anomaly] represent a powerful class of deep probabilistic generative models widely applied across various practical domains. These models incorporate a latent variable $\mathbf{z}$, sampled from prior distribution $p\left(\mathbf{z}\right)$, and conditional distribution $p_{\theta}\left(\mathbf{x}|\mathbf{z},\mathbf{c}\right)$ with respect to conditional label $\mathbf{c}$ to generate the observed variable $\mathbf{x}$. However, obtaining the marginal likelihood $p_{\theta}\left(\mathbf{x}\right)$ directly, becomes computationally intractable in high-dimensional settings due to the integration over the latent space. To address this challenge, variational inference is employed to approximate the posterior distribution and derive a tractable lower bound on the log-likelihood of the observed data. This results in the well-known evidence lower bound (ELBO), which serves as the objective function for training the model:
@@
Here, $q_{\phi}\left(\mathbf{z}|\mathbf{x},\mathbf{c}\right)$ denotes the variational approximation to the true posterior distribution $p_{\theta}\left(\mathbf{z}\mid\mathbf{x}\right)$, where $\phi$ and $\theta$ represent the parameters of the encoder and decoder networks, respectively. These components are typically implemented using neural networks. The CVAE is trained by minimizing the variational objective $\mathcal{L}\left(\mathbf{x};\theta,\phi\right)$ over the training data. However, the CVAE does not directly maximize the true likelihood; instead, it optimizes the ELBO, as defined in (7). This approximation can lead to learned representations that are overly constrained or limited in expressiveness, as only a limited subset of samples are likely to fall within regions of high posterior probability under the approximate posterior. To address this limitation, the importance-weighted autoencoder (IWAE) [burda2015importance] is employed. The IWAE introduces a tighter lower bound on the marginal log-likelihood $\log p\left(\mathbf{x}\right)$ by using a $k-$ sample importance-weighted estimate, thereby enabling a more expressive generative model.
@@ V-A Numerical Implementation and Training Results
The dataset used in this study is sourced from [git_UMG], as utilized in [hassler2023cyber]. All data were normalized using the min-max scaling approach as described in [hassler2023cyber]. After preprocessing, the final cyber-physical dataset comprises 12,514 instances, each with 30 features. This dataset was partitioned into training and testing subsets using an 80-20 split. To establish a baseline IDS, a multi-class feedforward neural network classifier was implemented. The network consists of three hidden layers, each containing 128 neurons. Its convergence behavior over training epochs is illustrated in Figure 3a. The trained IDS achieved a perfect classification accuracy of 100% on the test set.
@@ V-C Stealthy Attack Detection
In this subsection, we evaluate the effectiveness of different detection metrics in distinguishing stealthy generative adversarial attack from out-of-distribution (OOD) sample. As illustrated in Figure 6, negative log-likelihood (NLL) offers superior discriminatory power between adversarial and OOD samples compared to alternative metrics such as likelihood regret [xiao2020likelihood] and Mahalanobis distance [pang2018towards]. While likelihood regret is capable of differentiating benign inputs from OOD samples to some extent, it fails to account for the varying characteristics of distinct OOD inputs. This limitation arises because likelihood regret relies on an relative scoring mechanism, where the reference likelihood can become unstable under various nature of OOD distribution. As a result, its sensitivity to nuanced variations in OOD distributions is reduced. Similarly, the Mahalanobis distance assumes that features within each class are drawn from well-behaved Gaussian distributions. However, this assumption does not hold in the case of CVAE, whose learned latent representations often deviate from strict Gaussianity. Empirically, this observation is supported by the area under the curve (AUC) scores, where NLL achieves a value close to 0.99 as shown in Fig. 7. This indicates that NLL provides a more robust and reliable signal for distinguishing both OOD and adversarial inputs, outperforming the other two metrics in this setting.

### v2

# Generative Adversarial Evasion and Out-of-Distribution Detection for UAV Cyber-Attacks

@@ I INTRODUCTION
Unmanned Aerial Vehicles (UAVs) are increasingly employed in civilian applications, including aerial surveillance, precision agriculture, and logistics [10]. As critical components of emerging urban air mobility (UAM) systems, UAVs rely on networked communication and onboard autonomy, rendering them vulnerable to sophisticated cyber-attacks [7]. Adversaries may intercept telemetry data and inject falsified information by gaining unauthorized access to communication channels, thereby obstructing the UAM operator’s ability to verify compliance with planned flight paths as shown in Fig. 1. Failing to detect adversarial attacks as general anomalies can allow malicious actors to bypass security measures and inflict serious harm on airspace operations. This may lead to UAVs breaching restricted or unauthorized zones, thereby endangering other airspace users and heightening the risk of hijacking or unintentional exposure of sensitive information. Therefore, enhancing the security of UAV systems is crucial—especially in dynamic, time-critical operational environments. Traditional security methods such as cryptography and anomaly-based IDS face challenges like latency, high false positives, and poor generalization to novel threats [1].
@@
Generative adversarial networks (GANs) have emerged as a powerful tool for crafting polymorphic cyberattacks that evade detection. Prior work has demonstrated their use in generating adversarial examples that mimic legitimate data distributions while compromising classifier performance [3, 4]. However, distinguishing such stealthy adversarial samples from natural out-of-distribution (OOD) samples remains a major challenge. Adversaries can exploit this weakness by generating perturbations that resemble benign anomalies, thereby bypassing anomaly detection-based IDS. An inability to differentiate out-of-distribution (OOD) samples from adversarial attacks can undermine the effectiveness of conventional filtering methods, such as the extended Kalman filter (EKF) [14], which may struggle to detect stealthy adversarial inputs due to their statistical resemblance to the assumed OOD features.
Statistical methods for adversarial detection [8, 11] often assume fixed data distributions and lack adaptability to high-dimensional stealthy attacks. Moreover, they fail to model vulnerabilities in the latent space, limiting their effectiveness in detecting unseen or well-crafted adversarial inputs. To overcome these limitations, deep generative models such as variational autoencoders (VAEs) have been proposed [12]. VAEs capture probabilistic latent representations that support better anomaly detection. However, their generalization ability is limited in the presence of multiple attack types. Conditional VAEs (CVAEs) [9], which incorporate attack labels into the learning process, enhance discriminative power and robustness—but their application in detecting and characterizing diverse stealthy attacks remains underexplored. To address these gaps, this paper introduces a novel framework that combines generative adversarial evasion with probabilistic detection using CVAE for UAV cyber-physical systems. The key contributions are summarized as:
@@
  • Probabilistic Detection via CVAE: To detect such stealthy attacks from OOD samples, we implement a CVAE-based detection system. Using negative log-likelihood as a detection score, our method outperforms traditional Mahalanobis-based detectors [8] and advanced regret-based CVAE [15] baselines in distinguishing stealthy attacks from genuine OOD samples.
@@ II-B Cyber Attack Dataset
The UAV cyber-attack dataset utilized in this study is based on the intrusion detection framework developed by Hassler et al. [5], where a comprehensive experimental setup was employed. The dataset is comprised of UAV telemetry obtained after several types of cyberattacks, briefly described as follows:
@@
These attacks exhibit clear manifestations across both cyber and physical domains. Regarding the cyber features, only those attributes common across all attack types were retained to ensure consistency. These include: timestamp, frame number, frame length, transmission duration over the wireless medium, frame sequence number, control type and subtype, and the WLAN fragment number. In contrast, the physical features encompass various UAV telemetry data, such as altitude; velocities along the $x,y,z$ axes pitch, roll, yaw; flight distance and time; pressure; battery status; and spatial distances along the $x,y,z$ axes relative to the Tello Pad. Details of the experiment is provided in [5].
@@ III-B Discriminator
During the cGAN training, we assume that the attacker has the access to the IDS via an external API. The cGAN can be trained in a standard manner as shown in [16].
@@ III-C Iterative Refinement for Stealthy Attack
Iterative refinement strategy is analogous to the one in projected gradient descent (PGD) attacks [6], where attack is injected in multi-steps so that the overall perturbations stay within a certain bound. The aim is to make the adversarial attack $X_{\textup{adv}}$ close to the OOD sample $X_{\textup{ood}}$ generated from normally distributed noise with the noise scale $\varrho$, defined as $X_{\text{ood}}=X_{\text{att}}+\mathcal{N}\left(\mathbf{0},{\varrho}^{2}\mathbf{I}\right)$ . Hence, to ensure the stealth nature of the adversarial attack, $X_{\text{adv}}$ needs to be distributionally similar to $X_{\text{ood}}$, with high attack success rate. Hence, we define the stealthy adversarial attacks as follows:
@@ IV Conditional Variational Autoencoder for Stealthy Attack Detection
CVAEs [9] represent a powerful class of deep probabilistic generative models widely applied across various practical domains. These models incorporate a latent variable $\mathbf{z}$, sampled from prior distribution $p\left(\mathbf{z}\right)$, and conditional distribution $p_{\theta}\left(\mathbf{x}|\mathbf{z},\mathbf{c}\right)$ with respect to conditional label $\mathbf{c}$ to generate the observed variable $\mathbf{x}$. However, obtaining the marginal likelihood $p_{\theta}\left(\mathbf{x}\right)$ directly, becomes computationally intractable in high-dimensional settings due to the integration over the latent space. To address this challenge, variational inference is employed to approximate the posterior distribution and derive a tractable lower bound on the log-likelihood of the observed data. This results in the well-known evidence lower bound (ELBO), which serves as the objective function for training the model:
@@
Here, $q_{\phi}\left(\mathbf{z}|\mathbf{x},\mathbf{c}\right)$ denotes the variational approximation to the true posterior distribution $p_{\theta}\left(\mathbf{z}\mid\mathbf{x}\right)$, where $\phi$ and $\theta$ represent the parameters of the encoder and decoder networks, respectively. These components are typically implemented using neural networks. The CVAE is trained by minimizing the variational objective $\mathcal{L}\left(\mathbf{x};\theta,\phi\right)$ over the training data. However, the CVAE does not directly maximize the true likelihood; instead, it optimizes the ELBO, as defined in (7). This approximation can lead to learned representations that are overly constrained or limited in expressiveness, as only a limited subset of samples are likely to fall within regions of high posterior probability under the approximate posterior. To address this limitation, the importance-weighted autoencoder (IWAE) [2] is employed. The IWAE introduces a tighter lower bound on the marginal log-likelihood $\log p\left(\mathbf{x}\right)$ by using a $k$- sample importance-weighted estimate, thereby enabling a more expressive generative model.
@@ V-A Numerical Implementation and Training Results
The dataset used in this study is sourced from [13], as utilized in [5]. All data were normalized using the min-max scaling approach as described in [5]. After preprocessing, the final cyber-physical dataset comprises 12,514 instances, each with 30 features. This dataset was partitioned into training and testing subsets using an 80-20 split. To establish a baseline IDS, a multi-class feedforward neural network classifier was implemented. The network consists of three hidden layers, each containing 128 neurons. Its convergence behavior over training epochs is illustrated in Figure 3a. The trained IDS achieved a perfect classification accuracy of 100% on the test set.
@@ V-C Stealthy Attack Detection
In this subsection, we evaluate the effectiveness of different detection metrics in distinguishing stealthy generative adversarial attack from out-of-distribution (OOD) sample. As illustrated in Figure 6, negative log-likelihood (NLL) offers superior discriminatory power between adversarial and OOD samples compared to alternative metrics such as likelihood regret [15] and Mahalanobis distance [8]. While likelihood regret is capable of differentiating benign inputs from OOD samples to some extent, it fails to account for the varying characteristics of distinct OOD inputs. This limitation arises because likelihood regret relies on an relative scoring mechanism, where the reference likelihood can become unstable under various nature of OOD distribution. As a result, its sensitivity to nuanced variations in OOD distributions is reduced. Similarly, the Mahalanobis distance assumes that features within each class are drawn from well-behaved Gaussian distributions. However, this assumption does not hold in the case of CVAE, whose learned latent representations often deviate from strict Gaussianity. Empirically, this observation is supported by the area under the curve (AUC) scores, where NLL achieves a value close to 0.99 as shown in Fig. 7. This indicates that NLL provides a more robust and reliable signal for distinguishing both OOD and adversarial inputs, outperforming the other two metrics in this setting.